OpenStudy (anonymous):

State the horizontal asymptote of the rational function. f(x) = x + 9 / x^2 + 8x + 8

OpenStudy (debbieg):

There is a rule for horizonal asymptotes, related to the degree of the num'r vs. the degree of the den'r. Does that sound familiar?

OpenStudy (anonymous):

Yes, so it is horizontal because the denominator is bigger than the numerator

OpenStudy (anonymous):

One way of viewing it is to plug in a very large number x :) and evaluate

OpenStudy (debbieg):

Well, that isn't what makes it a HORIZONTAL asymptote. There are 2 cases where you get a horizontal asymyptote: deg num'r < deg den'r deg num'r = deg den'r

OpenStudy (anonymous):

(x + 9) / (x^2 + 8x + 8) just pick x= 999

OpenStudy (debbieg):

This is the first of those, so it gives a specific result. Think of it this way: for a BIG value of x, the den'r is going to be MUCH BIGGER than the num'r, so it will "take over". What will that do to the value of the rational expression?

OpenStudy (anonymous):

Or graphically https://www.google.com/search?q=(x+%2B+9)+%2F+(x%5E2+%2B+8x+%2B+8)&oq=(x+%2B+9)+%2F+(x%5E2+%2B+8x+%2B+8)&aqs=chrome..69i57.658j0&sourceid=chrome&ie=UTF-8 Just zoom out and scroll to the right and you will zee that the asymptote is... :D

OpenStudy (jdoe0001):

in the previous case the numerator's "degree" was bigger than the denominator's in this case is the other way around, notice the degree of the numerator is 1 and the denominator's degree is 2 so you do have a horizontal asymptote, what is it? well, is at y =0, or the x-axis

OpenStudy (anonymous):

x/x^2 here is another example the x^2 will get very big very fast as you aproach big numbers like 100 100/10000 = very small number 1000/1000000= even smaller number and you could say number is so small it aproaches 0

OpenStudy (anonymous):

so y = x

OpenStudy (jdoe0001):

well the x-axis line is y = 0

OpenStudy (anonymous):

I would say 0 :)

OpenStudy (anonymous):

here another example